OpenStudy (anonymous):

The Natchez Trace Bridge in Franklin, Tennessee, is 1,500 feet long. Suppose you build a model of the bridge using a scale of 1 inch:25 feet. How long is the model? (Hint: 12 inches = 1 foot). a. 5 ft c. 7 ft b. 6 ft d. 8 ft

OpenStudy (texaschic101):

1/25 = x/1500 -- ( 1 inch to 25 ft = x inch to 1500 ft) this is a proportion, so we will cross multiply 25x = (1)(1500) 25x = 1500 -- divide both sides by 25 x = 1500/25 x = 60 inches 60 inches = 5 ft (A)

OpenStudy (anonymous):

thank you

OpenStudy (texaschic101):

sure thing :)